OpenStudy (anonymous):

will Become Fan and give Metal!!! 3. Kathleen had a bag of candy-coated chocolates. Of the 60 candies in the bag, 20% were red, 30% were orange, 30% were green, and 20% were blue. Of the blue ones, 20% had a damaged candy coating. Approximately how many of the blue ones had a damaged candy coating? 24 2 12 6

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

@ertdre

OpenStudy (anonymous):

i need a few more correct and i pass the class

OpenStudy (anonymous):

How many blue candies does she have?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

ooo now i get it.

OpenStudy (anonymous):

And how many of those are damaged?

OpenStudy (anonymous):

you multyply 12 by 0.20 and get the Correct one. :p 2.4

OpenStudy (anonymous):

ty you helped me

OpenStudy (anonymous):

can i ask you a few more?

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

i have like 2 more or more to pass

OpenStudy (anonymous):

(3/3) 100.0% Overall Score: (42/62) 67.7%

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

4. Dave rented a limousine for his wife's birthday. The hourly rate is $80. They used the limousine for 5 hours, plus Dave gave the driver a 10% tip. How much did he spend in total for the hourly charges plus tip? $488 $480 $360 $440

OpenStudy (anonymous):

$440 ???

OpenStudy (anonymous):

How many hours did he use the limousine and what did it cost per hour....

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

80*5=400*0.10=40 400+40=440 ty

OpenStudy (anonymous):

i had my brother in colage help me he gave up :p

OpenStudy (anonymous):

last one 5. Brad had a small gathering at a local steakhouse. The steakhouse offers three dinner platters which vary by size and price. They ordered 4 of the 6-ounce platters, 3 of the 8-ounce platters, and 2 of the 11-ounce platters. Steak Platter Prices • 6-ounce $9.95 • 8-ounce $12.95 • 11-ounce $15.95 If a gratuity of 18% was added to the bill, which of the following is closest to the total of the steak platters and gratuity, ignoring sales tax? $154 $20 $130 $134

OpenStudy (anonymous):

Figure it like you did the last one...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

130.449? so c?

OpenStudy (anonymous):

Winner winner chicken dinner..great job

OpenStudy (anonymous):

;D ty i gave you a metal and added you as a fan.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

... i need 1 more

OpenStudy (anonymous):

K

OpenStudy (anonymous):

6. Sandy works at a clothing store. She makes $9 per hour plus earns 10% commission on her sales. She worked 80 hours over the last two weeks and had a total of $2,731 in sales before taxes. Which of the following is closest to how much she will earn in hourly wages and commission for those two weeks? $345 $273 $720 $993

OpenStudy (anonymous):

im at 68.8

OpenStudy (anonymous):

%

OpenStudy (anonymous):

Why not go for 100 you know how to do this.

OpenStudy (anonymous):

;p be cause i on Score: (5/5) 100.0% Overall Score: (44/64) 68.8%

OpenStudy (anonymous):

yah never have my brother help you.

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

i can never get a 100% now

OpenStudy (anonymous):

So waht do you think the answer is?

OpenStudy (anonymous):

9*0.10=0.9 9+0.9=9.9*80=792

OpenStudy (anonymous):

is what i did

OpenStudy (anonymous):

what about her sales commission?

OpenStudy (anonymous):

i added that that was the 0.9

OpenStudy (anonymous):

$720??? or $993???

OpenStudy (anonymous):

10% of 2731 is

OpenStudy (anonymous):

273.1

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

No it is the base pay which is hours worked x hourly wage plus the commission.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do i add the hours and the commission?

OpenStudy (anonymous):

to the bace

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

(Hours worked times $9.00 ) + commission = Total Pay

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

720+??

OpenStudy (anonymous):

or 720+273.1?

OpenStudy (anonymous):

if i do that i get D

OpenStudy (anonymous):

@nelsonjedi

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ty ty ty ty ty ty i am at 69.2 % now

OpenStudy (anonymous):

and the Final question is 7. Molly is having her birthday party at Pizza Party Palace. The costs involved are listed below. Pizza Party Palace requires a 15% deposit prior to the date of the event. If the deposit has been paid, how much is still owed for Molly's party? Party Prices • Food & Drink $250 • Game Tokens $100 • Other $70 $126 $357 $294 $63 @nelsonjedi and @amistre64

OpenStudy (anonymous):

$357

OpenStudy (anonymous):

. . . 8. Sam opened a money-market account that pays 4% simple interest. He started the account with $7,000 and made no further deposits. When he closed the account, he had earned $840 in interest. How long did he keep his account open? 4 years 5 years 3 years 6 years

OpenStudy (anonymous):

69.7%

OpenStudy (anonymous):

357 is correct

OpenStudy (anonymous):

so abut 3 years?

OpenStudy (anonymous):

it is 3 or 4 years

OpenStudy (anonymous):

you are the best nelsonjedi

OpenStudy (anonymous):

How much interest does he earn each year?

OpenStudy (anonymous):

4%

OpenStudy (anonymous):

Money I mean?

OpenStudy (anonymous):

or 280

OpenStudy (anonymous):

$$

OpenStudy (anonymous):

omg @nelsonjedi I am So Happy about for you you got me a 70.1 % how can i ever repay you lol :p

OpenStudy (anonymous):

Thanks is enough..congrats

OpenStudy (anonymous):

ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty ty.

OpenStudy (anonymous):

now i am tire and going to bed Night Cya